OpenStudy (anonymous):

Fred is fencing his yard, which is shaped as a rhombus. He used 4x meters of fencing material for one side and (x + 9) meters for the adjacent side. What is the perimeter of the yard? 12 meters 36 meters 96 meters 48 meters

jimthompson5910 (jim_thompson5910):

all sides are congruent, so 4x = x+9 solve that for x to get x = _____

OpenStudy (anonymous):

the question its given he used 4x meters for fencing one side that means 4x = 4*3 =12m ( since x=3) and x+9 for other side (3)+9 =12m So you have one side of Rhombus as 12m Perimeter = Sum of all sides = 12+12+12+12 (or 4 * 12) =48 m